OpenStudy (anonymous):

If i pass 3 quarters and fail one quarter do i still pass?

OpenStudy (alexwee123):

depends what did you fail in? o.0

OpenStudy (anonymous):

well i passed all three quarters in geometry this quarter i failed

OpenStudy (mertsj):

(0+60+60+60)/4=180/4=45 Is 45 passing?

OpenStudy (alexwee123):

well geometry is required course for grad o,0

OpenStudy (alexwee123):

for high school

OpenStudy (anonymous):

merts

OpenStudy (anonymous):

what are you trying to say

OpenStudy (mertsj):

I'm trying to say that it depends on the percentages you received each quarter.

OpenStudy (anonymous):

ok can i tell you

OpenStudy (anonymous):

Q1:B Q2:C Q3:D Q4:F

OpenStudy (anonymous):

what were the percentages?

OpenStudy (mertsj):

Are you on a 4 point scale at your school?

OpenStudy (anonymous):

here wait

OpenStudy (anonymous):

doesnt say percent

OpenStudy (mertsj):

A=4, B=3, C=2, D=1, F=0 (3+2+1+0)?4=1.5=C

OpenStudy (anonymous):

so whats your saying is i failed

OpenStudy (mertsj):

Is C failing?

OpenStudy (anonymous):

o i got a c?

OpenStudy (mertsj):

Did you even read what I wrote?

OpenStudy (anonymous):

o sorry didnt see

OpenStudy (anonymous):

wait what about a,c,f,f

OpenStudy (mertsj):

A=4, B=3, C=2, D=1, F=0 (3+2+1+0)/4=1.5=C

OpenStudy (mertsj):

(4+1+0+0)/4=1.25=C-

OpenStudy (anonymous):

so i got a c overall?

OpenStudy (mertsj):

If I were the teacher, yes. However, I would give you a final exam and it would count 1/5 of the grade...the same as each quarter.

OpenStudy (anonymous):

ok so if i do good on my exam im coolio right?

OpenStudy (mertsj):

Has your teacher discussed how the grades are going to be calculated? I can't speak for your teacher.

OpenStudy (anonymous):

no clue man no clue

OpenStudy (anonymous):

i just wanna pass all my classes

OpenStudy (anonymous):

and get freshmen year done with all the classes

OpenStudy (mertsj):

I hope you do. Good luck.

OpenStudy (anonymous):

so my worst grades will be c's

OpenStudy (anonymous):

all my other grades will be b's and a's

OpenStudy (anonymous):

hell yeah

OpenStudy (mertsj):

Well hopefully you will do better next year. Remember: The more you know, the further you go.

OpenStudy (anonymous):

yeah that and i cant see the board D:

OpenStudy (mertsj):

Get some glasses.

OpenStudy (anonymous):

75% ?